# Chest X-ray Foundation Model with Global and Local Representations Integration

### v1

@@ 3.2 Global and Local Representation Integration for Multilabel Classification
where $\mathbf{q}^{\prime}\in\mathbb{R}^{M\times D_{\text{key}}}$ denotes the output disease queries, which we consider contain disease-specific local features. $D_{\text{key}}$ is the dimension of the key tokens. Last, we concatenate $\mathbf{q}^{\prime}$ with the [CLS] token to construct the GLoRI output token sequence. For multilabel classification, each GLoRI output token is projected by a linear classifier supervised by a binary cross-entropy loss.
Overall, GLoRI applies the attentional principle, which shares similarity with previous works, such as Perceiver [39] and DETR [40], for CheXFound’s downstream evaluation. We empirically demonstrate that GLoRI extracts local features relevant to disease abnormalities in CXRs in Section 6.1.
Figure 2: Global and Local Representations Integration (GLoRI) for evaluating CheXFound on downstream tasks. GLoRI is appended on top of the frozen CheXFound backbone. GLoRI uses disease queries to compute attention-pooled CXR representations and integrates the global image features from [CLS] token for disease finding classification. For the opportunistic CXR interpretation, additional disease queries are created for GLoRI to retrieve representations related to the CVD risk estimation and mortality prediction.
@@ 4.1.1 Self-supervised pretraining details
We conducted self-supervised pretraining of CheXFound on our curated CXR-1M dataset (Table 1) using the ViT-L architecture with a patch size of 16$\times$16 pixels. The loss weights for $\mathcal{L}_{\texttt{[CLS]}}$ and $\mathcal{L}_{\texttt{MIM}}$ were set to 1.0 and 3.0, respectively. The momentum to compute the exponential moving average of the student network was set to 0.994. We varied the global and local crop sizes to pretrain CheXFound at a rang of resolutions. Specifically, we set the global and local crop size pairs to be (512, 144), (448, 128), (336, 128), and (224, 96). We set the number of global and local crops to 2 and 8, respectively. For masked image modeling, we set the proportion of masked patches to the range (0.1, 0.5). We trained CheXFound for 100 epochs with an epoch length of 2,500 iterations and a batch size of 14 per graphics processing unit (GPU). We used the AdamW optimizer with an initial learning rate of 2e-4. We applied a Cosine annealing schedule for learning rate decay and a warm up period of 10 epochs. We pretrained CheXFound on a DGX-1 server with 8×8$\times$8 × NVIDIA A100 40GB GPUs. Depending on the image resolutions, the pretraining processes take around 48 to 96 hours. The project source code is publicly available at https://github.com/RPIDIAL/CheXFound.
@@ 4.1.2 Downstream evaluation details
In downstream evaluation, we trained GLoRI with feature representations from the frozen CheXFound. We took the concatenated representations from the last 4 layers of CheXFound as the input to GLoRI. We set the embedding dimension of GLoRI to 768. Disease queries were randomly initialized with a standard normal distribution. For the cross-attention layer, we used the multihead attention mechanism and set the number of heads to 8. To train GLoRI, we used the AdamW optimizer and conduct a thorough learning rate search in {1e-5, 2e-5, 5e-5, 1e-4, 2e-4, 5e-4, 1e-3, 2e-3, 5e-3} to obtain the best-performing learning rate on the validation set. We then combined the training and validation sets for a second round of training using the best learning rate. To maximize the number of images that the GLoRI module processes during downstream adaptation, we trained GLoRI for 10 epochs on CXR-LT 24, CheXpert, and PLCO and 100 epochs on Shenzhen, Mongomery, and JSRT. We set the batch size to 16 for Montgomery and JSRT and 256 for CXR-LT 24, CheXpert, Shenzhen, and PLCO.
4.2 Experimental Design
To rigorously evaluate CheXFound’s in-distribution and out-of-distribution performance, we employed an extensive classification benchmark, consisting of in-distribution datasets (CXR-LT 24 [22] and CheXpert [15]) and out-of-distribution datasets (Shenzhen [41], Montgomery[41], Japanese Society of Radiological Technology (JSRT) [42], and Prostate, Lung, Colorectal, and Ovarian (PLCO) Cancer Screening Trial [23]). Following recent works in vision-centric foundation models [12, 26], we split the evaluation datasets into training, validation, and test splits. To avoid any potential data contamination, we included only the training set in CXR-LT 24 and the training and validation sets in CheXpert for self-supervised pretraining, while keeping the test set unseen. For the out-of-distribution datasets (Shenzhen, Montgomery, and JSRT), none of the images in the training, validation, and test sets were used for self-supervised pretraining.
@@ 4.2 Experimental Design
To assess the out-of-distribution generalization capabilities of the foundation models, we also performed evaluation on the Shenzhen and Montgomery datasets for tuberculosis detection and the JSRT database for lung nodule detection. The Shenzhen, Montgomery, and JSRT datasets were divided into training, validation, and test splits with a ratio of 70:10:20. Shenzhen contains 463 training images, 65 validation images, and 134 test images. Montgomery contains 96 training images, 14 validation images, and 28 test images. JSRT contains 171 training images, 24 validation images, and 50 test images.
@@
Since the CXR interpretation problem often involves severe class imbalance, we employed two metrics to evaluate model performance: the area under the precision-recall curve (AUPRC) and the area under the receiver operating characteristic curve (AUROC). For the multilabel classification problem, we computed the average of the metrics over disease findings. We estimated the 95% confidence intervals of the model performance in AUPRC and AUROC over 1,000 bootstrapped samples. To test statistical significance, we used a two-sided paired permutation test with 1,000 permutations to assess the observed performance differences of the two models for disease findings.
@@ 5 Experimental Results
Table 2: Comparison CheXFound and other foundation models when using linear probe and GLoRI for classification. Values inside the parentheses indicate the 95% confidence intervals. Values in bold indicate the best results.
Classifier Foundation CXR-LT 24 CheXpert Shenzhen Montgomery JSRT
@@
Linear probe PubMedCLIP [29] 0.089(.088–.089) 0.561(.554–.568) 0.277(.244–.311) 0.595(.564–.628) 0.857(.786–.913) 0.814(.738–.887) 0.565(.306–.817) 0.534(.310–.750) 0.785(.630–.912) 0.685(.511–.833)
BiomedCLIP [28] 0.117(.116–.118) 0.643(.636–.649) 0.557(.504–.607) 0.841(.822–.860) 0.903(.843–.949) 0.885(.827–.934) 0.925(.791–1.0) 0.929(.822–1.0) 0.589(.427–.755) 0.432(.265–.601)
CheXzero [13] 0.112(.111–.112) 0.552(.545–.558) 0.468(.424–.509) 0.778(.754–.798) 0.929(.884–.962) 0.906(.851–.950) 0.964(.869–1.0) 0.970(.898–1.0) 0.759(.597–.885) 0.620(.453–.777)
EVA-X [11] 0.114(.113–.115) 0.596(.590–.602) 0.468(.422–.512) 0.788(.763–.812) 0.840(.738–.921) 0.824(.746–.898) 0.577(.320–.809) 0.509(.278–.749) 0.641(.472–.797) 0.490(.312–.660)
RAD-DINO [12] 0.114(.113–.114) 0.557(.570–.583) 0.463(.422–.503) 0.746(.715–.778) 0.883(.818–.933) 0.861(.798–.916) 0.637(.369–.842) 0.561(.316–.788) 0.747(.585–.888) 0.623(.456–.778)
CheXFound 0.209(.204–.214) 0.799(.794–.804) 0.620(.630–.727) 0.876(.860–.892) 0.974(.949–.992) 0.967(.935–.990) 0.988(.939–1.0) 0.990(.952–1.0) 0.918(.826–.975) 0.856(.741–.948)
GLoRI head PubMedCLIP [29] 0.116(.115–.117) 0.649(.643–.655) 0.501(.450–.552) 0.804(.778–.828) 0.897(.832–.946) 0.867(.807–.927) 0.628(.363–.866) 0.694(.484–.872) 0.652(.483–.810) 0.493(.330–.654)
BiomedCLIP [28] 0.122(.121–.123) 0.643(.636–.649) 0.552(.506–.593) 0.829(.809–.847) 0.921(.870–.957) 0.897(.840–.944) 0.900(.738–1.0) 0.898(.744–1.0) 0.634(.470–.789) 0.505(.291–.704)
CheXzero [13] 0.131(.130–.132) 0.671(.665–.677) 0.599(.551–.647) 0.888(.868–.905) 0.912(.852–.954) 0.894(.838–.942) 0.726(.451–.901) 0.653(.396–.882) 0.640(.467–.807) 0.485(.323–.662)
EVA-X [11] 0.149(.147–.150) 0.679(.672–.685) 0.614(.571–.659) 0.870(.853–.888) 0.928(.881–.964) 0.896(.835–.945) 0.977(.909–1.0) 0.979(.918–1.0) 0.866(.760–.936) 0.748(.615–.864)
RAD-DINO [12] 0.173(.171–.176) 0.723(.717–.729) 0.639(.593–.687) 0.884(.869–.898) 0.909(.854–.952) 0.885(.823–.936) 0.909(.747–1.0) 0.911(.782–1.0) 0.683(.503–.861) 0.615(.440–.783)
CheXFound 0.252(.247–.258) 0.830(.826–.834) 0.679(.630–.727) 0.908(.894–.921) 0.983(.960–.996) 0.978(.951–.995) 1.000(1.0–1.0) 1.000(1.0–1.0) 0.986(.956–1.0) 0.975(.931–1.0)
@@
GLoRI head CheXFound 0.252(.247–.258) 0.830(.826–.834) 0.679(.630–.727) 0.908(.894–.921) 0.983(.960–.996) 0.978(.951–.995) 1.000(1.0–1.0) 1.000(1.0–1.0) 0.986(.956–1.0) 0.975(.931–1.0)
@@ 5.1 Overall performance comparison
A pivotal characteristic of foundation models lies in their capability to achieve improved performance on a wide range of downstream datasets. To evaluate the capability of foundation models, we compared CheXFound, which uses ViT-L pretrained on CXR-1M, with publicly available pretrained encoders, including RAD-DINO [12], EVA-X [11], CheXzero [13], BiomedCLIP [28], PubmedCLIP [29], and ConvNeXt [43]. RAD-DINO was pretrained on a combined dataset comprising MIMIC-CXR, CheXpert, PadChest, CXR14, and BRAX, using the DINOv2 framework. EVA-X was pretrained on MIMIC-CXR, CheXpert, and CXR14 using EVA [44] technique with contrastive vision features for masked image modeling. CheXzero, a vision-language foundation model, was pretrained on MIMIC-CXR using contrastive language-image pretraining (CLIP) [45]. BiomedCLIP was pretrained on PMC-15M with image-text pairs collected from scientific articles. PubmedCLIP was initialized with the CLIP model and finetuned on PubMed articles. Finally, ConvNeXt was pretraiend on the ImageNet-22K dataset.
@@
To evaluate the effectiveness and generalizability of the representations extracted by the foundation models, we evaluated the linear probe performance across the five datasets (CXR-LT 24, CheXpert, Shenzhen, Montgomery, and JSRT). While linear probe provides a straightforward approach to evaluate the quality of representation, it primarily relies on global image features from the [CLS] token, often resulting in suboptimal performance. Hence, we also evaluated the pretrained encoders using GLoRI, which incorporates attention-pooled local features in addition to the global image features from the [CLS] token. For vision-language pretained encoders, we further validated the quality of their vision representations by examing their correlation with text features for disease finding classification.
@@
Across all the five classification tasks (CXR-LT 24, CheXpert, Shenzhen, Montgomery, and JSRT), CheXFound consistently outperformed other foundation models in the linear probe setting as shown in Table 2. On the multilabel, long-tailed classification task (CXT-LT 24), CheXFound achieved an AUPRC of 0.209, outperforming the next best-performing model (either RAD-DINO or EVA-X) by 9.5% ($p<0.001$, two-sided paired permutation test). On the five-class multilabel classification task (CheXpert), CheXFound outperformed the next best-performing model (EVA-X) by 8.8% ($p<0.001$) in AUROC. On single-class classification tasks (Shenzhen, Montgomery, and JSRT) with limited amounts of training data, CheXFound similarly outperformed the next best-performing models in AUROC by 11.7% ($p<0.001$), 42.9% ($p<0.001$), and 23.3% ($p<0.001$), respectively.
@@
We further evaluated the performance of the foundation models using GLoRI across five classification tasks (Table 2). The foundation models (CheXFound, RAD-DINO, EVA-X, CheXzero, BiomedCLIP, and PubMedCLIP) with GLoRI generally outperformed their linear probe baselines. Specifically, CheXFound with GLoRI outperformed its linear probe baseline in AUROC by 3.1%, 3.2%, 1.1%, 1.0%, and 11.9% on CXR-LT 24, CheXpert, Shenzhen, Montgomery, and JSRT, respectively. CheXFound with GLoRI also outperformed other foundation models, including CLIP-based models (CheXzero, BioMedCLIP, and PubMedCLIP) under image-text alignment, as well as end-to-end trained ConvNeXt (Table 3).
@@
To show the detailed performance of CheXFound over 40 disease findings, we illustrated CheXFound performance in AUROC against its comparisons in Fig. 3. CheXFound consistently outperformed other methods across the 40 disease findings in AUROC. We also compared performance in AUPRC over disease findings with high, medium, and low prevalence in Fig. 3b. CheXFound outperformed its comparisons under all levels of prevalence, even for underrepresented pathologies in the low prevalence category.

### v2

@@ 3.2.3 Integrating Global Image Features
Last, we integrate the fine- and coarse-grained local features with the [CLS] token to construct the GLoRI output token sequence. For multilabel classification, each GLoRI output token represents a disease finding and is projected by a linear classifier, optimized using a binary cross-entropy loss, to perform disease detection.
Overall, GLoRI employs the attention pooling mechanism, a strategy also employed by CLIP [49], V-JEPA [50], RayDINO [34], and GLoRIA [51] that incorporated attention pooling for image representation pooling [49, 51], video classification [50], and disease detection [34]. GLoRI distinguishes itself from these related works by integrating two local-feature branches to capture both fine- and coarse-grained local features. We empirically demonstrate that GLoRI can extract local features relevant to disease abnormalities in CXRs in Section 6.1.
Figure 2: Global and Local Representations Integration (GLoRI) for evaluating CheXFound on downstream tasks. GLoRI is trained on top of the frozen CheXFound backbone. GLoRI incorporates adaptive temperatures and pyramid patch merging into the attention mechanism to encode fine-grained and coarse-grained local features respectively and integrates the global image features from the [CLS] token to predict final multilabel classification results.
@@ 4.1.1 Self-supervised pretraining details
We conducted self-supervised pretraining of CheXFound on our curated CXR-987K dataset (Table 2) using the ViT-L architecture with a patch size of 16$\times$16 pixels. The loss weights for $\mathcal{L}_{\texttt{[CLS]}}$ and $\mathcal{L}_{\texttt{MIM}}$ were set to 1.0 and 3.0, respectively. The momentum to compute the exponential moving average of the student network was set to 0.994. We varied the global and local crop sizes to pretrain CheXFound at a range of resolutions. Specifically, we set the global and local crop size pairs to be (512, 144), (448, 128), (336, 128), and (224, 96). We set the number of global and local crops to 2 and 8, respectively. For masked image modeling, we set the proportion of masked patches to the range (0.1, 0.5). We trained CheXFound for 100 epochs with an epoch length of 2,500 iterations and a batch size of 14 per graphics processing unit (GPU). We used the AdamW optimizer with an initial learning rate of 2e-4. We applied a Cosine annealing schedule for learning rate decay and a warm up period of 10 epochs. We pretrained CheXFound on a DGX-1 server with 8×8$\times$8 × NVIDIA A100 40GB GPUs. Depending on the image resolutions, the pretraining processes take around 48 to 96 hours.
@@ 4.1.2 Downstream evaluation details
In downstream evaluation, we trained GLoRI with feature representations from the frozen CheXFound. We took the concatenated representations from the last 4 layers of CheXFound as the input to GLoRI. We set the embedding dimension of GLoRI to 768. Disease queries were randomly initialized with a standard normal distribution. For the cross-attention layer, we used the multihead attention mechanism and set the number of heads to 8. To train GLoRI, we used the AdamW optimizer and conduct a thorough learning rate search in {1e-5, 2e-5, 5e-5, 1e-4, 2e-4, 5e-4, 1e-3, 2e-3, 5e-3} to obtain the best-performing learning rate on the validation set. We then combined the training and validation sets for a second round of training using the best learning rate. To maximize the number of images that the GLoRI head processes during downstream adaptation, we trained GLoRI for 10 epochs on CXR-LT 24, CheXpert, and PLCO and 100 epochs on Shenzhen, Mongomery, and JSRT. We set the batch size to 16 for Montgomery and JSRT and 256 for CXR-LT 24, CheXpert, Shenzhen, and PLCO.
4.2 Experimental datasets
To rigorously evaluate CheXFound’s in-distribution and out-of-distribution performance, we employed an extensive classification benchmark, consisting of in-distribution datasets (CXR-LT 24 [21] and CheXpert [15]) and out-of-distribution datasets (VinDr-CXR [23], CXR-Pneumonia [52], Shenzhen [53], Montgomery[53], Japanese Society of Radiological Technology (JSRT) [54], RANZCR-CLiP [24], and Prostate, Lung, Colorectal, and Ovarian (PLCO) Cancer Screening Trial [22]). Following recent works in vision-centric foundation models [12, 28], we split the evaluation datasets into training, validation, and test splits. To avoid any potential data contamination, we included only the training set in CXR-LT 24 and the training and validation sets in CheXpert for self-supervised pretraining, while keeping the test set unseen. For the out-of-distribution datasets (VinDr-CXR, CXR-Pneumonia, Shenzhen, Montgomery, JSRT, RANZCR-CLiP, and PLCO), none of the images in the training, validation, and test sets were used for self-supervised pretraining.
@@ 4.2 Experimental datasets
To assess the out-of-distribution generalization capabilities of the foundation models, we evaluated model performance on VinDr-CXR that contains labels for 28 diseases. We used the official data splits of VinDr-CXR that contain 15000 CXRs for training and 3000 CXRs for test. We also performed evaluation on CXR-Pneumonia for pneumonia detection, Shenzhen and Montgomery for tuberculosis detection, and JSRT for lung nodule detection. The CXR-Pneumonia dataset contains 5,232 / 624 CXRs for training and test respectively. The Shenzhen, Montgomery, and JSRT datasets were divided into training, validation, and test splits with a ratio of 70:10:20. Shenzhen contains 463 training images, 65 validation images, and 134 test images. Montgomery contains 96 training images, 14 validation images, and 28 test images. JSRT contains 171 training images, 24 validation images, and 50 test images. In addition, we evaluated model performance for malpositioned tube detection on RANZCR-CLiP. We split the RANZCR-CLiP dataset into 18,742 / 2,678 / 5,356 CXRs for training, validation, and test respectively.
@@
Since the CXR interpretation problem often involves severe class imbalance, we employed two metrics to evaluate model performance: the area under the precision-recall curve (AUPRC) and the area under the receiver operating characteristic curve (AUROC). For the multilabel classification problem, we computed the average of the metrics over disease findings. We estimated the 95% confidence intervals of the model performance in AUPRC and AUROC over 1,000 bootstrapped samples. To test statistical significance, we used a two-sided permutation test with 1,000 permutations to assess the observed performance differences of the two models for disease findings.
@@ 5 Experimental Results
Table 3: Classification results of CheXFound and other foundation models using linear probe and GLoRI on in-distribution datasets. Values inside the parentheses indicate the 95% confidence intervals. Values in bold indicate the best results.
Classifier Foundation CXR-LT 24 CheXpert
Methods Models AUPRC AUROC AUPRC AUROC
Linear probe PubMedCLIP [39] 0.089(.088–.089) 0.561(.554–.568) 0.277(.244–.311) 0.595(.564–.628)
BiomedCLIP [38] 0.117(.116–.118) 0.643(.636–.649) 0.557(.504–.607) 0.841(.822–.860)
CheXzero [13] 0.112(.111–.112) 0.552(.545–.558) 0.468(.424–.509) 0.778(.754–.798)
EVA-X [11] 0.114(.113–.115) 0.596(.590–.602) 0.468(.422–.512) 0.788(.763–.812)
RAD-DINO [12] 0.114(.113–.114) 0.557(.570–.583) 0.463(.422–.503) 0.746(.715–.778)
CheXFound 0.209(.204–.214) 0.799(.794–.804) 0.620(.630–.727) 0.876(.860–.892)
GLoRI head PubMedCLIP [39] 0.116(.115–.117) 0.649(.643–.655) 0.501(.450–.552) 0.804(.778–.828)
BiomedCLIP [38] 0.122(.121–.123) 0.643(.636–.649) 0.552(.506–.593) 0.829(.809–.847)
CheXzero [13] 0.131(.130–.132) 0.671(.665–.677) 0.599(.551–.647) 0.888(.868–.905)
EVA-X [11] 0.149(.147–.150) 0.679(.672–.685) 0.614(.571–.659) 0.870(.853–.888)
RAD-DINO [12] 0.173(.171–.176) 0.723(.717–.729) 0.639(.593–.687) 0.884(.869–.898)
CheXFound 0.265(.259–.271) 0.840(.836–.844) 0.679(.630–.727) 0.908(.894–.921)
Table 4: Classification results of CheXFound and other foundation models on out-of-distribution datasets. Values inside the parentheses indicate the 95% confidence intervals. Values in bold indicate the best results.
Classifier Foundation VinDr-CXR CXR-Pneumonia Shenzhen Montgomery JSRT
@@
Linear probe PubMedCLIP [39] 0.088(.082–.095) 0.570(.514–.609) 0.954(.936–.969) 0.924(.904–.944) 0.857(.786–.913) 0.814(.738–.887) 0.565(.306–.817) 0.534(.310–.750) 0.785(.630–.912) 0.685(.511–.833)
BiomedCLIP [38] 0.100(.094–.106) 0.587(.544–.618) 0.984(.976–.990) 0.973(.962–.986) 0.903(.843–.949) 0.885(.827–.934) 0.925(.791–1.0) 0.929(.822–1.0) 0.589(.427–.755) 0.432(.265–.601)
CheXzero [13] 0.102(.096–.109) 0.536(.486–.567) 0.973(.962–.981) 0.953(.938–.967) 0.929(.884–.962) 0.906(.851–.950) 0.964(.869–1.0) 0.970(.898–1.0) 0.759(.597–.885) 0.620(.453–.777)
EVA-X [11] 0.105(.099–.111) 0.542(.503–.572) 0.975(.959–.988) 0.968(.953–.980) 0.840(.738–.921) 0.824(.746–.898) 0.577(.320–.809) 0.509(.278–.749) 0.641(.472–.797) 0.490(.312–.660)
RAD-DINO [12] 0.102(.096–.107) 0.477(.441–.507) 0.995(.991–.998) 0.992(.986–.996) 0.883(.818–.933) 0.861(.798–.916) 0.637(.369–.842) 0.561(.316–.788) 0.747(.585–.888) 0.623(.456–.778)
CheXFound 0.298(.280–.317) 0.838(.776–.866) 0.996(.993–.998) 0.993(.988–.997) 0.974(.949–.992) 0.967(.935–.990) 0.988(.939–1.0) 0.990(.952–1.0) 0.918(.826–.975) 0.856(.741–.948)
GLoRI head PubMedCLIP [39] 0.104(.098–.113) 0.589(.532–.640) 0.973(.961–.983) 0.960(.944–.973) 0.897(.832–.946) 0.867(.807–.927) 0.628(.363–.866) 0.694(.484–.872) 0.652(.483–.810) 0.493(.330–.654)
BiomedCLIP [38] 0.094(.089–.102) 0.537(.497–.566) 0.987(.981–.992) 0.978(.969–.986) 0.921(.870–.957) 0.897(.840–.944) 0.900(.738–1.0) 0.898(.744–1.0) 0.634(.470–.789) 0.505(.291–.704)
CheXzero [13] 0.074(.072–.078) 0.489(.446–.515) 0.982(.972–.989) 0.972(.959–.982) 0.912(.852–.954) 0.894(.838–.942) 0.726(.451–.901) 0.653(.396–.882) 0.640(.467–.807) 0.485(.323–.662)
EVA-X [11] 0.163(.155–.172) 0.662(.610–.694) 0.990(.979–.997) 0.988(.978–.995) 0.928(.881–.964) 0.896(.835–.945) 0.977(.909–1.0) 0.979(.918–1.0) 0.866(.760–.936) 0.748(.615–.864)
RAD-DINO [12] 0.155(.147–.163) 0.680(.630–.718) 0.994(.991–.997) 0.991(.985–.995) 0.909(.854–.952) 0.885(.823–.936) 0.909(.747–1.0) 0.911(.782–1.0) 0.683(.503–.861) 0.615(.440–.783)
CheXFound 0.341(.319–.367) 0.872(.813–.898) 0.996(.992–.998) 0.993(.987–.997) 0.983(.960–.996) 0.978(.951–.995) 1.000(1.0–1.0) 1.000(1.0–1.0) 0.986(.956–1.0) 0.975(.931–1.0)
@@
GLoRI head CheXFound 0.265(.259–.271) 0.840(.836–.844) 0.679(.630–.727) 0.908(.894–.921) 0.983(.960–.996) 0.978(.951–.995) 1.000(1.0–1.0) 1.000(1.0–1.0) 0.986(.956–1.0) 0.975(.931–1.0)
@@ 5.1 Overall performance comparison
A pivotal characteristic of foundation models lies in their capability to achieve improved performance on a wide range of downstream datasets. To evaluate the capability of foundation models, we compared CheXFound, which uses ViT-L pretrained on CXR-987K, with publicly available pretrained encoders, including RAD-DINO [12], EVA-X [11], CheXzero [13], BiomedCLIP [38], PubmedCLIP [39], and ConvNeXt [55]. RAD-DINO was pretrained on a combined dataset comprising MIMIC-CXR, CheXpert, PadChest, NIH-CXR, and BRAX, using the DINOv2 framework. EVA-X was pretrained on MIMIC-CXR, CheXpert, and NIH-CXR using EVA [56] technique with contrastive vision features for masked image modeling. CheXzero, a vision-language foundation model, was pretrained on MIMIC-CXR using contrastive language-image pretraining (CLIP) [49]. BiomedCLIP was pretrained on PMC-15M with image-text pairs collected from scientific articles. PubmedCLIP was initialized with the CLIP model and finetuned on PubMed articles. Finally, ConvNeXt was pretrained on the ImageNet-22K dataset.
@@
To evaluate the effectiveness and generalizability of the representations extracted by the foundation models, we evaluated the linear probe performance across seven datasets (CXR-LT 24, CheXpert, VinDr-CXR, CXR-Pneumonia, Shenzhen, Montgomery, and JSRT). While linear probe stands for a straightforward approach to evaluate the quality of representations, it primarily relies on global image features from the [CLS] token, often resulting in suboptimal performance. We hence evaluated the pretrained encoders using our proposed GLoRI, which incorporates attention-pooled local features in addition to the global image features to enhance multilabel classification. Regarding vision-language pretrained encoders, we further validated their vision representation quality by assessing their correlation with text features to perform zero-shot classification.
@@
Across all the in-distribution and out-of-distribution classification tasks (CXR-LT 24, CheXpert, VinDr-CXR, CXR-Pneumonia, Shenzhen, Montgomery, and JSRT), CheXFound consistently outperformed other foundation models in the setting using a simple linear probe classifier, as shown in Tables 3 and 4. On the multilabel, long-tailed classification task (CXT-LT 24), CheXFound achieved an AUPRC of 0.209, outperforming the next best-performing model (either RAD-DINO or EVA-X) by 9.5% ($p<0.001$, two-sided permutation test). On the five-class multilabel classification task (CheXpert), CheXFound outperformed the next best-performing model (BiomedCLIP) by 3.5% ($p<0.001$) in AUROC. On the out-of-distribution multilabel classification task (VinDr-CXR), CheXFound outperformed the next best-performing model (BiomedCLIP) by 25.1% ($p<0.001$) in AUROC. On out-of-distribution single-class classification tasks (CXR-Pneumonia, Shenzhen, Montgomery, and JSRT) with limited amounts of training data, CheXFound outperformed the next best-performing models in AUROC by 0.1% ($p>0.05$), 6.1% ($p<0.001$), 2.0% ($p>0.05$), and 17.1% ($p<0.001$), respectively.
@@
We further evaluated the performance of the foundation models using GLoRI across the seven classification tasks (Tables 3 and 4). The foundation models (CheXFound, RAD-DINO, EVA-X, CheXzero, BiomedCLIP, and PubMedCLIP) with GLoRI generally outperformed their linear probe baselines. CheXFound with GLoRI outperformed its linear probe baseline in AUROC by 4.1%, 3.2%, 3.4%, 1.1%, 1.0%, and 11.9% on CXR-LT 24, CheXpert, VinDr-CXR, Shenzhen, Montgomery, and JSRT, respectively. CheXFound achieved an AUROC of 0.993 on CXR-Pneumonia, on par with its linear probe baseline. In addition, as shown in Table 5, CheXFound with GLoRI outperformed CLIP-based foundation models (CheXzero, BioMedCLIP, and PubMedCLIP) when image-text alignment was used for zero-shot classification. CheXFound also outperformed the end-to-end trained model ConvNeXt.
@@
To show the detailed performance of CheXFound over 40 disease findings, we illustrated CheXFound performance in AUROC against its comparisons in Fig. 3. CheXFound consistently outperformed other methods across the 40 disease findings in AUROC. We also compared performance in AUPRC over disease findings with high, medium, and low prevalence in Fig. 4. CheXFound outperformed its comparisons under all levels of prevalence, even for underrepresented pathologies in the low prevalence category.
5.2 Malpositioned tube detection
The RANZCX-CLiP dataset contains CXRs for verifying the position of lines and tubes. This dataset categorized these CXRs into normal, borderline, and abnormal classes based on the position of tubes (endotracheal tube, nasogastric tube, and central venous catheter) to indicate whether they were appropriately positioned. The percentages of normal, borderline, and abnormal samples are 59.4%, 28.8%, and 11.8% respectively, representing an unbalanced class distribution. As shown in Table 6, CheXFound achieved AUPRCs of 0.857, 0.529, and 0.445 for classifying normal, borderline, and abnormal tubes respectively. CheXFound outperformed the next best-performing models by 16.4% ($p<0.001$), 16.1% ($p<0.001$), and 29.2% ($p<0.001$) for classifying normal, borderline, and abnormal classes respectively. These performance gains could be attributed to CheXFound’s capability to encode high-resolution CXRs and produce effective representations of nuanced structures.
Table 6: Classification results of foundation models for malpositioned tube detection in CXRs. We computed the mean AUPRC values and their 95% confidence intervals over 1,000 bootstrapped samples. Values in bold indicate the best-performing results.
Methods Normal Borderline Abnormal
PubMedCLIP [39] 0.661(.642–.679) 0.348(.328–.372) 0.136(.123–.149)
BiomedCLIP [38] 0.645(.626–.664) 0.330(.311–.350) 0.133(.121–.148)
CheXzero [13] 0.693(.674–.711) 0.364(.341–.386) 0.153(.139–.173)
EVA-X [20] 0.657(.637–.676) 0.339(.317–.361) 0.135(.122–.150)
RAD-DINO [20] 0.690(.670–.709) 0.368(.347–.391) 0.138(.125–.154)
CheXFound 0.857(.844–.869) 0.529(.503–.556) 0.445(.403–.488)